OpenStudy (anonymous):

find max and min value of f(x) = 2x^2-8x+1 by completing the square. show work please

OpenStudy (anonymous):

completing squares I don't know that method I will like to see how it's done

OpenStudy (anonymous):

f(x)=2(x^2-4x+4)-7 =2(x-2)^2-7 so (x-2)^2 is always +ve so f(x)min = -7 and f(x) max is infinity

OpenStudy (anonymous):

f(x) has no maximum.

OpenStudy (anonymous):

your statement is partially correct @pol f(x) has no finite maximum thats why i wrote f(x) max is to be infinite